Help please
11 of 40 (0 comp his Question: 1 pt The ages, in years, of 6 people who applied for a job are: 21, 40, 23, 25, 37, 34 What is the midrange age? OA, 28 ○B.30.5 OC. 22.5 ○ D, correct answer is not given |
Answer #1

Maximum: Max = 40

Minimum: Min = 21

The midrange:

Midrange=\frac{Max+Min}{2}=\frac{40+21}{2}=30.5

Correct option is B.